OpenStudy (dschneider2016):

Will medal and fan!! Just check my work? Verify the identity. cos 4x + cos 2x = 2 - 2 sin^2 2x - 2 sin^2 x

OpenStudy (dschneider2016):

@hartnn @ParthKohli Please help!! I really need it

OpenStudy (dschneider2016):

@poopsiedoodle Please could you just check my work?

OpenStudy (dschneider2016):

cos 4x +cos 2x = 2 cos ^2 2x -1 + 2 cos^2 x -1 = 2(1 - sin^2 2x ) -1 +cos 2x =1-2sin^2 2x + 1-2sin^2 x =2 - 2sin^2 2x - 2 sin^2 x

OpenStudy (dschneider2016):

@raffle_snaffle

OpenStudy (anonymous):

Nicely done!

OpenStudy (anonymous):

Let me check one more time.

OpenStudy (dschneider2016):

are you sure its right because Im really not definite?

OpenStudy (anonymous):

Yes, you are very much correct!

OpenStudy (raffle_snaffle):

great he helped you

OpenStudy (raffle_snaffle):

=)

OpenStudy (anonymous):

That's offensive @raffle_snaffle

OpenStudy (raffle_snaffle):

i guess it came out wrong. lol I am glad you helped him

OpenStudy (dschneider2016):

okay so I shouldn't include anything else in my response to be more thorough? @eggs&bacon

OpenStudy (dschneider2016):

her*

OpenStudy (anonymous):

Nope, I think it reads pretty well.

OpenStudy (raffle_snaffle):

lol sorry her...

OpenStudy (anonymous):

@raffle_snaffle Lol, why I said I'm offended is because I'm a girl! Lol

OpenStudy (raffle_snaffle):

oh lol sorry

OpenStudy (dschneider2016):

@gorv

OpenStudy (anonymous):

@dsneider2016 How'd you know I was a girl?

OpenStudy (anonymous):

Oh wait, you're a girl. Nevermind.

OpenStudy (anonymous):

I know, I know. Lol. Did I help?

OpenStudy (gorv):

yes your method is correct bud

OpenStudy (dschneider2016):

Oh thank you!

OpenStudy (dschneider2016):

so you as well wouldnt add anything?

OpenStudy (gorv):

nope

OpenStudy (gorv):

you did it in right way

OpenStudy (dschneider2016):

ok Thank you!

OpenStudy (anonymous):

Wait, why he get medal and me don't?